# UNI: Unlearning-based Neural Interpretations

### v1

@@ 1 Introduction
Gradient-based saliency methods are widely used for feature attribution, due to their simplicity, efficiency and post-hoc accessibility. This can be further decomposed into 3 families: perturbative, backpropagative and path-based, which we detail in Section 6. Gradient-based attribution is intuitive since the first-order derivative reveals which features significantly influence the model’s classification decision. However, naively using local gradients yields unfaithful attributions due to saturation, where the non-linear output function flattens in vicinity of the input and zero gradients are computed (Sundararajan et al., 2017; 2016). To improve gradient-sensitivity, later methods introduce a baseline input for reference, and backpropagate the difference in activation scores on a path between the reference and image-of-interest (Shrikumar et al., 2016; Sundararajan et al., 2017). The baseline is chosen to be devoid of predictive features and far away from the saturated local neighbourhood. However, such methods accumulate gradient noise when interpolating from the baseline to the input, leading to high local sensitivity (Ancona et al., 2018). Consequently, attribution maps become disconnected, sparse and irregular, where the saliency scores fluctuate wildly between neighbouring pixels of the same object and are visually noisy (Adebayo et al., 2018). This noise accumulation has two root causes—a poorly chosen baseline and high-curvature output manifold along the path features. Previous works (Sturmfels et al., 2020; Xu et al., 2020) have sought better baselines by empirically comparing between using a black image, a gaussian noised image, a gaussian blurred image, a uniformly noised image, an inverted colour image, as well as averaging attributions over several baseline choices. However, the correct baseline to represent a lack of salient features depends heavily on the specific classification task, on the trained model and on the input image. Indeed, the optimal baseline varies for each task–model–image combination (Akhtar & Jalwana, 2023); the baseline problem remains largely unsolved. Turning to the second problem of high-curvature output manifold, because trained neural networks exhibit approximately piece-wise linear decision boundaries (Goodfellow et al., 2014), therefore inputs near function transitions are vulnerable to perturbative attacks. By simply adding norm-bounded, imperceptible adversarial noise to the input image, attackers can dramatically alter the attribution map without changing the model’s class prediction (Ghorbani et al., 2019a; Dombrowski et al., 2019). Methods of mitigation include explicit smoothing via averaging over multiple noised gradient attributions (Smilkov et al., 2017); adaptively optimising the integration path of attribution (Kapishnikov et al., 2021); imposing an attribution prior during training and optimising it at each step (Erion et al., 2021). However, all of these proposals starkly increase the complexity of attribution, requiring computationally costly forward and backward propagation steps.
@@
    Post-hoc attribution can impose new biases. We approach the baseline challenge from the fresh lens of post-hoc biases. We show that static baselines (e.g. black, blurred, random noise) inject additional colour, texture and frequency assumptions that are not present in the original model’s decision rule, which leads explanation infidelity and inconsistency.
@@
    Unlearning reduces the curvature of decision boundaries and increases robustness. Targeted unlearning simulates the function statistics of unseen data, and smooths the curvature of the output manifold around the sample. This is characterised by low geodesic curvature of the path and bounded principal curvature of the output surface. This points to reduced variability of the gradient vector under small-norm input perturbations, leading to better attribution robustness and faithfulness.
@@ 3 Gradient-based Attributions in a Nutshell
2:Choose unlearning step-size $\eta$;           PGD steps $T$, budget $\epsilon$, step-size $\mu$;       Riemann approximation steps $B$
@@ 3.2 Post-hoc Biases are Imposed
Since the baseline represents an absence of or reduction in salient features, static baseline functions (e.g. black, blurred, noised) implicitly assume that similar features (e.g. dark, smooth, high-frequency) are irrelevant for model prediction. To illustrate this intuition, we can consider IG with a black baseline, wherein it becomes more difficult to attribute dark but salient pixels. Due to the colour bias that “near-black features are unimportant", the term $(x_{i}-x^{\prime}_{i})$ is small and requires a disproportionately large gradient $\nabla_{x_{i}}F_{c}(\cdot)$ to yield non-negligible attribution scores. Indeed, this is what we observe in Figures 2, 5, 8, where darker features belonging to the object-of-interest cannot be reliably identified. We further empirically verify that each static baseline imposes its own post-hoc bias by experimenting on ImageNet-C (Hendrycks & Dietterich, 2019). Corresponding to the 3 popular baseline choices for IG (all-black, gaussian blurred, gaussian noised), we focus on the families of digital (brightening and saturation), blur (gaussian and defocus blur) and noise (gaussian and shot noise) common-corruptions. Figures 5, 9 demonstrate that IG with a blurred baseline fails to attribute blurred inputs due to saturation and overly smoothed image textures; Figures 5, 10 visualise how a noised IG baseline encounters high-frequency noise and outputs irregular, high-variance attribution scores, even for adjacent pixels belonging to the same object. We crucially emphasise that such colour, texture and frequency biases are not present naturally in the pre-trained model but rather injected implicitly by a suboptimal choice of static baseline. The observation that poor baseline choices create attribution bias has so far been overlooked. As such, we depart entirely from the line of work on alternative static baseline towards adaptively (un)learning baselines with gradient-based optimisation. UNI eliminates all external assumptions except for the model’s own predictive bias.
@@ 4.1 Baseline desiderata
A desirable baseline should preserve the game-theoretic properties of path-attribution (Section 3.1) and refrain from imposing post-hoc attribution biases (Section 3.2). For every given task-model-image triad, a well-chosen baseline should be 1. image-specific—be connected via a path feature of low curvature to the original image; 2. reflect only the model’s predictive biases—salient image features should be excluded from the baseline; be 3. less task-informative than the original image—interpolating from the baseline towards the input image should yield a path of increasing predictive confidence. We now introduce the UNI pipeline: first, unlearn predictive information in the model space; then, use activation-matching between unlearned and trained models to mine a featurelesss baseline in the image space; finally, interpolate along the low-curvature, conformant and consistent path from baseline to image to compute reliable explanations in the attributions space.
@@ 4.2 Desirable Path Features
The meaningfulness of the attributions highly depends on the meaningfulness of the path. We aim for a smooth transition between absence and presence of features; and this intuitively cannot be achieved if the baseline and input are too far apart. (Srinivas & Fleuret, 2019) formalizes this intuition through the concept of weak dependence, and proves that this property can only be compatible with completeness in the case where the baseline and the input lie in the same connected component (in the case of piecewise-linear models). An obvious implementation of this proximity condition in the general case is to bound the distance $||x-x^{\prime}||$ to a certain value $\varepsilon$. This is strictly enforced in Algorithm 1 by normalizing the perturbation at each step $t$.
@@
The curvature of the model prediction along the integrated path has been identified (Dombrowski et al., 2019) as one of the key factors influencing both the sensitivity and faithfulness of the computed attributions. We substantiate the intuition that a smooth and regular path is preferred by analysing the Riemannian sum calculation. Assuming that the function $g:\alpha\in[0,1]\mapsto\nabla F_{c}\left(x^{\prime}+\alpha(x-x^{\prime})\right)$ is derivable with a continuous derivative (i.e. $\mathcal{C}^{1}$) on the segment $[x^{\prime},x]$, elementary calculations and the application of the Taylor-Lagrange inequality give the following error in the Riemann approximation of the attribution,
@@
Thus, lower curvature along the path implies a lower value of the constant $M$, which in turn implies a lower error in the integration calculation. A smaller value $B$ of Riemann steps is needed to achieve the same precision. More generally, a low curvature (i.e. eigenvalues of the hessian) on and in a neighborhood of the baseline and path reduces the variability of the calculated gradients under small-norm perturbations, increasing the sensitivity and consistency of the method. Empirically, we observe a much lower curvature of the paths computed by UNI, as can be seen in Table LABEL:tab:mono and Appendix Figures 17, 18, 19, 20, 21, 22. Figure 7 also confirms the increased robustness to Riemann sum error induced.
@@
Intuitively, the path $\gamma$ defined by interpolating from the “featureless" baseline $x^{\prime}$ to the input image $x$ should be monotonically increasing in output class confidence; the cumulative attribution score should similarly increase. At the image level, for all $j,k$ such that $j\leq k$, since $|\gamma(j)-x|\geq|\gamma(k)-x|$, therefore the predictive confidence should be non-decreasing and order-preserving: $F_{c}(\gamma(j))\leq F_{c}(\gamma(k))$. Constraining $\gamma$ to be monotonically increasing suffices to satisfy criteria for valid path features (Akhtar & Jalwana, 2023). Zooming in on input feature $\tilde{x}_{i}=\gamma(\tilde{k})_{i}$ on the linear path for $\tilde{k}<1$, since the gradient of $\gamma$ is always positive and $x_{i}$ maximises $F_{c}$ on the interval, therefore both conditions $\textrm{sgn}(\nabla_{x_{i}}F_{c}(x_{i}))\cdot\textrm{sgn}(\nabla_{\tilde{x}_{i}}F_{c}(\tilde{x}_{i}))=1$ and $|\nabla_{x_{i}}F_{c}(x_{i})|>|\nabla_{\tilde{x}_{i}}F_{c}(\tilde{x}_{i})|$ are naturally met. This is crucial for preserving the completeness axiom of Integrated Gradients (Sundararajan et al., 2017), that attribution scores add up to the difference quantity $\sum_{i=1}^{d_{\mathcal{X}}}\mathcal{A}_{i}(x)=F_{c}(x)-F_{c}(x^{\prime})$, and that attribution score $\mathcal{A}_{i}(x)$ is only non-zero when $x_{i}$ contributes to the prediction. Completeness does not hold when path features are non-conformant, i.e. $F_{c}(\gamma(j))>F_{c}(\gamma(k))$ ; or when extremities exist along $\gamma$, i.e. for $\tilde{k}\in(0,1),\>F_{c}(\gamma(\tilde{k}))>F_{c}(\gamma(1))$ or $F_{c}(\gamma(\tilde{k}))<F_{c}(\gamma(0))$.
@@ 5 Experiments
We experiment on ImageNet-1K (Deng et al., 2009), ImageNet-C (Hendrycks & Dietterich, 2019) and compare against various path-based and gradient-based attribution methods. This includes IG (Sundararajan et al., 2017), BlurIG (Xu et al., 2020), GIG (Kapishnikov et al., 2021), AGI (Pan et al., 2021), GBP (Springenberg et al., 2014) and DeepLIFT (Shrikumar et al., 2016). We consider a diverse set of pre-trained computer vision backbone models (Paszke et al., 2019), including ResNet-18 (He et al., 2016), EfficientNet-v2-small (Tan & Le, 2021), ConvNeXt-Tiny (Liu et al., 2022), VGG-16-bn (Simonyan & Zisserman, 2015), ViT-B_16 (Dosovitskiy et al., 2020) and Swin-Transformer-Tiny (Liu et al., 2021). Unless otherwise specified, we the following hyperparameters: unlearning step size $\eta=1$; $l_{2}$ PGD with $T=10$ steps, a budget of $\varepsilon=0.25$, step size $\mu=0.1$; Riemann approximation with $B=15$ steps. Our results verify UNI’s high faithfulness, stability and robustness.
5.1 Faithfulness
We report MuFidelity scores (Bhatt et al., 2021), i.e. the faithfulness of an attribution function $\mathcal{A}$, to a model $F$, at a sample $x$, for a subset of features of size $|S|$, given by $\mu_{f}(F,\mathcal{A};x)=\textrm{corr}_{S\in{[d]\choose|S|}}\left(\sum_{i\in S}\mathcal{A}(i,F,c,x),\>F_{c}(x)-F_{c}(x_{[x_{s}=\bar{x}_{s}]})\right)$. We record the (absolute) correlation coefficient between a randomly sampled subset of pixels and their attribution scores. As from Table 2, UNI outperforms other methods across all settings but one, indicating high faithfulness. We supplement these numbers with visual comparisons in Appendix Figures 11, 12, 13, 14, 15, 16 against IG (black and noised baselines), BlurIG, GIG, AGI, GBP, DeepLift.
Table 2: MuFidelity scores measure the correlation between a subset of pixels’ impact on the output (i.e. change in predictive confidence) and assigned saliency scores. Since attribution methods can yield strong positive or negative correlation, we report the absolute scores.

### v2

# Unlearning-based Neural Interpretations

@@ 1 Introduction
Gradient-based saliency methods are widely used for feature attribution, due to their simplicity, efficiency and post-hoc accessibility. This can be further decomposed into 3 families: perturbative, backpropagative and path-based, which we detail in Section 6. Gradient-based attribution is intuitive since the first-order derivative reveals which features significantly influence the model’s classification decision. However, naively using local gradients yields unfaithful attributions due to saturation, where the non-linear output function flattens in vicinity of the input and zero gradients are computed (Sundararajan et al., 2017; 2016). To improve gradient-sensitivity, later methods introduce a baseline input for reference, and backpropagate the difference in activation scores on a path between the reference and image-of-interest (Shrikumar et al., 2016; Sundararajan et al., 2017). The baseline is chosen to be devoid of predictive features and far away from the saturated local neighbourhood. However, such methods accumulate gradient noise when interpolating from the baseline to the input, leading to high local sensitivity (Ancona et al., 2018). Consequently, attribution maps become disconnected, sparse and irregular, where the saliency scores fluctuate wildly between neighbouring pixels of the same object and are visually noisy (Adebayo et al., 2018). This noise accumulation has two root causes—a poorly chosen baseline and high-curvature output manifold along the path features. Previous works (Sturmfels et al., 2020; Xu et al., 2020) have sought better baselines by empirically comparing between using a black image, a gaussian noised image, a gaussian blurred image, a uniformly noised image, an inverted colour image, as well as averaging attributions over several baseline choices. However, the correct baseline to represent a lack of salient features depends heavily on the specific classification task, on the trained model and on the input image. Indeed, the optimal baseline varies for each task–model–image combination (Akhtar & Jalwana, 2023); the baseline problem remains largely unsolved. Turning to the second problem of high-curvature output manifold, because trained neural networks exhibit approximately piece-wise linear decision boundaries (Goodfellow et al., 2014), inputs near function transitions are vulnerable to perturbative attacks. By simply adding norm-bounded, imperceptible adversarial noise to the input image, attackers can dramatically alter the attribution map without changing the model’s class prediction (Ghorbani et al., 2019a; Dombrowski et al., 2019). Methods of mitigation include explicit smoothing via averaging over multiple noised gradient attributions (Smilkov et al., 2017); adaptively optimising the integration path of attribution (Kapishnikov et al., 2021); imposing an attribution prior during training and optimising it at each step (Erion et al., 2021). However, all of these proposals starkly increase the complexity of attribution, requiring computationally costly forward and backward propagation steps.
@@
    Post-hoc attribution can impose new biases. We approach the baseline challenge from the fresh lens of post-hoc biases. We show that static baselines (e.g. black, blurred, random noise) inject additional colour, texture and frequency assumptions that are not present in the original model’s decision rule, which leads to explanation infidelity and inconsistency.
@@
    Unlearning reduces the curvature of decision boundaries and increases robustness. Targeted unlearning simulates function statistics of unseen data, and smooths the curvature of the output manifold around the sample. This is characterised by low geodesic path curvature and bounded principal curvature of the output surface. This points to reduced variability of the gradient vector under small-norm input perturbations, leading to better attribution robustness and faithfulness.
@@ 3 Gradient-based Attributions in a Nutshell
2:Choose unlearning step-size $\eta$; PGD steps $T$,         budget $\varepsilon$, step-size $\mu$; Riemann approximation steps $B$
@@ 3.2 Post-hoc Biases are Imposed
Since the baseline represents an absence of or reduction in salient features, static baseline functions (e.g. black, blurred, noised) implicitly assume that similar features (e.g. dark, smooth, high-frequency) are irrelevant for model prediction. To illustrate this intuition, we can consider IG with a black baseline, wherein it becomes more difficult to attribute dark but salient pixels. Due to the colour bias that “near-black features are unimportant", the term $(x_{i}-x^{\prime}_{i})$ is small and requires a disproportionately large gradient $\nabla_{x_{i}}F_{c}(\cdot)$ to yield non-negligible attribution scores. Indeed, this is what we observe in Figures 2, 5, 11, where darker features belonging to the object-of-interest cannot be reliably identified. We further empirically verify that each static baseline imposes its own post-hoc bias by experimenting on ImageNet-C (Hendrycks & Dietterich, 2019). Corresponding to the 3 popular baseline choices for IG (all-black, gaussian blurred, gaussian noised), we focus on the families of digital (brightening and saturation), blur (gaussian and defocus blur) and noise (gaussian and shot noise) common-corruptions. Figures 5, 12 demonstrate that IG with a blurred baseline fails to attribute blurred inputs due to saturation and overly smoothed image textures; Figures 5, 13 visualise how a noised IG baseline encounters high-frequency noise and outputs irregular, high-variance attribution scores, even for adjacent pixels belonging to the same object. We crucially emphasise that such colour, texture and frequency biases are not present naturally in the pre-trained model but rather injected implicitly by a suboptimal choice of static baseline. The observation that poor baseline choices create attribution bias has so far been overlooked. As such, we depart entirely from the line of work on alternative static baseline towards adaptively (un)learning baselines with gradient-based optimisation. UNI eliminates all external assumptions except for the model’s own predictive bias.
@@ 4.1 Baseline desiderata
A desirable baseline should preserve the game-theoretic properties of path-attribution (Section 3.1) and refrain from imposing post-hoc attribution biases (Section 3.2). For every given task-model-image triad, a well-chosen baseline should be 1. image-specific—be connected via a path feature of low curvature to the original image; 2. reflect only the model’s predictive biases—salient image features should be excluded from the baseline; be 3. less task-informative than the original image—interpolating from the baseline towards the input image should yield a path of increasing predictive confidence. We now introduce the UNI pipeline: first, unlearn predictive information in the model space; then, use activation-matching between unlearned and trained models to mine a featureless baseline in the image space; finally, interpolate along the low-curvature, conformant and consistent path from baseline to image to compute reliable explanations in the attributions space. Figure 8 visuals and Table 8 results attest to UNI’s ability to compute specific, unlearned baselines for attribution.
@@ 4.2 Desirable Path Features
The meaningfulness of the attributions highly depends on the meaningfulness of the path. We aim for a smooth transition between absence and presence of features; and this intuitively cannot be achieved if the baseline and input are too far apart. Srinivas & Fleuret (2019) formalises this intuition through the concept of weak dependence, and proves that this property can only be compatible with completeness in the case where the baseline and the input lie in the same connected component (in the case of piecewise-linear models). An obvious implementation of this proximity condition in the general case is to bound the distance $||x-x^{\prime}||$ to a certain value $\varepsilon$. This is strictly enforced in Algorithm 1 by normalising the perturbation at each step $t$.
@@
The curvature of the model prediction along the integrated path has been identified Dombrowski et al. (2019) as one of the key factors influencing both the sensitivity and faithfulness of the computed attributions. We substantiate the intuition that a smooth and regular path is preferred by analysing the Riemannian sum calculation. Assuming that the function $g:\alpha\in[0,1]\mapsto\nabla F_{c}\left(x^{\prime}+\alpha(x-x^{\prime})\right)$ is derivable with a continuous derivative (i.e. $\mathcal{C}^{1}$) on the segment $[x^{\prime},x]$, elementary calculations and the application of the Taylor-Lagrange inequality give the following error in the Riemann approximation of the attribution,
@@
Thus, lower curvature along the path implies a lower value of the constant $M$, which in turn implies a lower error in the integration calculation. A smaller value $B$ of Riemann steps is needed to achieve the same precision. More generally, a low curvature (i.e. eigenvalues of the hessian) on and in a neighbourhood of the baseline and path reduces the variability of the calculated gradients under small-norm perturbations, increasing the sensitivity and consistency of the method. Empirically, we observe a much lower curvature of paths computed by UNI, as per Table LABEL:tab:mono and Appendix Figures 20, 21, 22, 23, 24, 25. Figure 10 also confirms the increased robustness to Riemann sum error induced.
@@
Intuitively, the path $\gamma$ defined by interpolating from the “featureless" baseline $x^{\prime}$ to the input image $x$ should be monotonically increasing in output class confidence. At the image level, for all $j,k$ such that $j\leq k$, since $||\gamma(j)-x||\geq||\gamma(k)-x||$, therefore the predictive confidence should be non-decreasing and order-preserving: $F_{c}(\gamma(j))\leq F_{c}(\gamma(k))$. Constraining $\gamma$ to be monotonically increasing suffices to satisfy a weak version of the criteria for valid path features (Akhtar & Jalwana, 2023): $\textrm{sgn}(\nabla_{x}F_{c}(x))\cdot\textrm{sgn}(\nabla_{\tilde{x}}F_{c}(x^{\prime}))=1$ is naturally met.
4.3 Effects of Unlearning and Matching
We explain the success of UNI with the illustrative example of a three gaussians mixture model. Figure 6 computes unlearning and activation matching for a model learned on three data points with gradient descent. $F$ is chosen to be the output of the three gaussian components $(G_{1},G_{2},G_{3})$. Note that the perturbation is not $\varepsilon$-normalised for clearer visualisation. We highlight two observations:
  • The UNI path is monotonous, of low-curvature and proximal. Conversely, the path to the random baseline is long, non-monotonous, and goes through several zones of high second derivative.
  • Optimizing KL divergence on $(G_{1},G_{2},G_{3})$ produces a better baseline. Figure 6(b) visualises the unlearning objective (i.e. the target probability after unlearning), which gives four points of intersection with the base model ($a,b,c$ and $d$). By constraining proximity of the baseline with the $\varepsilon$ parameter, we restrain the optima found by gradient descent (on the global probability) to the closest two points $a$ and $b$. UNI is then able yield the more optimal of the two, by optimising on each gaussian output. In fact, the idea of activation matching is to satisfy the crucial weak dependence property for conformal path attribution (Akhtar & Jalwana, 2023). Since modern ReLU networks have decision boundaries representable as piecewise linear functions (Xiong et al., 2020), activation matching supervises the baseline to use the same (piecewise linear) weights. In our case, we want to find a baseline for which $G_{1}$ and $G_{2}$ do not play a role, which is not the case for $a$. This is why Algorithm 1 optimises on $F$ and not on $F_{c}$.
Finally, $\varepsilon$ normalisation serves to regularise baseline GD learning and account for pathological cases where the locally shortest path would lead to further intersections than the closest one.
(a) Unlearned Model
(b) Base Model
Figure 6: UNI baseline on a Gaussian mixture model of three gaussians $G_{1},G_{2},G_{3}$, each of fixed variance, parametrised by their mean and a scaling factor. (b) shows the model trained on the three datapoints (1,1), (2, 5) and (4, 3), while (a) shows the model after one gradient ascent step on the datapoint (2, 5). The path between UNI Baseline and the image is highlighted by arrows in (b).
@@ 5 Experiments
We experiment on ImageNet-1K (Deng et al., 2009), ImageNet-C (Hendrycks & Dietterich, 2019) and compare against various path-based and gradient-based attribution methods. This includes IG (Sundararajan et al., 2017), BlurIG (Xu et al., 2020), GIG (Kapishnikov et al., 2021), AGI (Pan et al., 2021), GBP (Springenberg et al., 2014) and DeepLIFT (Shrikumar et al., 2016). We consider a diverse set of pre-trained computer vision backbone models (Paszke et al., 2019), including ResNet-18 (He et al., 2016), EfficientNet-v2-small (Tan & Le, 2021), ConvNeXt-Tiny (Liu et al., 2022), VGG-16-bn (Simonyan & Zisserman, 2015), ViT-B_16 (Dosovitskiy et al., 2020) and Swin-Transformer-Tiny (Liu et al., 2021). Unless otherwise specified, we the following hyperparameters: unlearning step size $\eta=1$; $l_{2}$ PGD with $T=10$ steps, a budget of $\varepsilon=0.25$, step size $\mu=0.1$; Riemann approximation with $B=15$ steps. We further extend UNI to the NLP domain, to interpret generative language models using activation patching (Heimersheim & Nanda, 2024). UNI complements activation matching by computing a stable baseline without trading off attribution scalability, as in Appendix Table 8 and Figure 9. Our results verify UNI’s high faithfulness, stability and robustness.
Table 2: MuFidelity scores measure the correlation between a subset of pixels’ impact on the output (i.e. change in predictive confidence) and assigned saliency scores. Since attribution methods can yield strong positive or negative correlations, we report the absolute scores.
@@
5.1 Faithfulness
We report MuFidelity scores (Bhatt et al., 2021), i.e. the faithfulness of an attribution function $\mathcal{A}$, to a model $F$, at a sample $x$, for a subset of features of size $|S|$, given by $\mu_{f}(F,\mathcal{A};x)=\textrm{corr}_{S\in{[d]\choose|S|}}\left(\sum_{i\in S}\mathcal{A}(i,F,c,x),\>F_{c}(x)-F_{c}(x_{[x_{s}=\bar{x}_{s}]})\right)$. We record the (absolute) correlation coefficient between a randomly sampled subset of pixels and their attribution scores. In line with open source exemplars (Fel et al., 2022a), we set $|S|$ to be 25% of the total pixel count (slightly higher than the referenced 20%) as is required to adjust for ImageNet’s complexity and for obtaining less noisy measurements across all baseline methods. As from Table 2, UNI outperforms other methods across all settings but one, indicating high faithfulness. We supplement these numbers with visual comparisons in Appendix Figures 14, 15, 16, 17, 18, 19 against IG (black and noised baselines), BlurIG, GIG, AGI, GBP, DeepLift. Furthermore, we report deletion and insertion scores (Petsiuk et al., 2018)—a causally-motivated evaluation metric for interpretability methods—which measures the decrease (deletion) or increase (insertion) of a model’s output confidence as salient pixels are removed (from the original image) or inserted (into a featureless baseline). A steep drop in model confidence under pixel deletion results in a desirable and small area under the curve (AUC) score; a sharp rise under pixel insertion results in a desirably large AUC. Salient pixels are removed in descending order of importance, as identified by the tested interpretability method. We evaluate with a step size of 10% and average over 10,000 random image samples, where at each step, the next-10% most salient pixels are removed or inserted for inference. UNI reliably identifies pixels which are crucial for sample classification, achieving marked improvements especially in insertion AUC scores.
Table 3: Deletion AUC $\downarrow$ measures how confidence drops as pixels are removed (lower = better).
UNI IG BlurIG GIG AGI GBP DeepLIFT
ResNet-18 .$\bm{.06}$.$\pm.128$$10$$\pm.174$$27$$\pm.252$11.11.11.11 ±.$\pm.150$$13$$\pm.147$$08$$\pm.160$$13$$\pm.165$
EfficientNetv2s .$19$$\pm.212$$26$$\pm.217$$50$$\pm.158$$19$$\pm.216$$\bm{.18}$.$\pm.207$$23$$\pm.163$$27$$\pm.215$
ConvNeXt-Tiny .$\bm{.11}$.$\pm.139$$16$$\pm.164$$46$$\pm.172$$21$$\pm.160$$17$$\pm.123$$16$$\pm.099$$21$$\pm.162$
VGG-16-bn .$\bm{.08}$.$\pm.143$$12$$\pm.181$$18$$\pm.241$$10$$\pm.163$$14$$\pm.178$$14$$\pm.194$$12$$\pm.186$
ViT-B_16 .$14$$\pm.185$22.22.22.22 ±.$\pm.207$$60$$\pm.166$$17$$\pm.190$$\bm{.13}$.$\pm.152$$23$$\pm.141$$17$$\pm.189$
Swin-T-Tiny .$\bm{.13}$.$\pm.181$22.22.22.22 ±.$\pm.217$$47$$\pm.174$22.22.22.22 ±.$\pm.207$$21$$\pm.172$$21$$\pm.123$$23$$\pm.207$
Table 4: Insertion AUC $\uparrow$ measures how confidence rises as pixels are inserted (higher = better).
UNI IG BlurIG GIG AGI GBP DeepLIFT
ResNet-18 .$\bm{.64}$.$\pm.138$$26$$\pm.045$$34$$\pm.131$$36$$\pm.048$$56$$\pm.068$$11$$\pm.066$$18$$\pm.042$
EfficientNetv2s .$\bm{.64}$.$\pm.227$$38$$\pm.127$$51$$\pm.283$$37$$\pm.138$$38$$\pm.204$$23$$\pm.192$$37$$\pm.137$
ConvNeXt-Tiny .$\bm{.63}$.$\pm.231$$21$$\pm.114$$40$$\pm.252$$56$$\pm.122$$52$$\pm.088$$22$$\pm.160$$17$$\pm.162$
VGG-16-bn .$\bm{.56}$.$\pm.335$$37$$\pm.061$$31$$\pm.274$$38$$\pm.071$$47$$\pm.078$$26$$\pm.057$$17$$\pm.056$
ViT-B_16 .$\bm{.71}$.$\pm.237$$32$$\pm.107$$59$$\pm.292$$28$$\pm.125$$43$$\pm.089$$35$$\pm.172$$28$$\pm.123$
Swin-T-Tiny .$\bm{.68}$.$\pm.245$$28$$\pm.145$$63$$\pm.282$$26$$\pm.153$$25$$\pm.156$$31$$\pm.202$$26$$\pm.152$